OpenStudy (anonymous):

an object is released from rest on a planet that has no atmosphere. the object falls freely for 5.82 m on the first second. what is the magnitude of the acceleration due to gravity on the planet

OpenStudy (anonymous):

11.64m/s^2 given..u=0m/s, n=1sec, S(n)=5.82m, a=+g(since, downward motion) from S(n)=u+a(n-0.5); 5.82=0+g(1-0.5) 5.82=g(0.5) g=5.82/0.5 g=11.64m/s^2 hope this helps you...cheers!!!